# Design and Verification of the JUNO Liquid Filling Control System

### v1

@@ 5 Conclusion
The FOC control system serves as the central platform for LS filling, overflowing, and circulating in the JUNO experiment. By integrating a high-stability PLC architecture, precision sensors (accuracy < 0.2% FS), and advanced control logic with PID algorithms, sequential control and safety interlocks, the system achieves closed-loop regulation of critical parameters with exceptional reliability. Test performances demonstrate the water/LS filling and the shock-free LS transfer with rapid stabilization in 3 minutes and flow rate control accuracy of 0.5%. Hardware redundancy such as modular PLCs, one-operational-one-standby pumps and software safeguards (e.g., dual HMIs, multi-type alarm mechanisms) ensure fault tolerance and anti-interference capability. These achievements establish a solid foundation for the safe and efficient filling of the 20-kiloton LS detector and its stable operation over 20-year lifespan.

### v3

@@ 5 Conclusion
The FOC system serves as the central platform for LS filling, overflow, and circulation in the JUNO experiment. Integrating a high-stability PLC architecture, precision sensors (accuracy < 0.2% FS), and advanced control strategies—including PID regulation, sequential logic, and safety interlocks—the system achieves closed-loop control of critical parameters with high reliability. Commissioning tests demonstrate rapid flow stabilization within 3 minutes and a control accuracy of 0.5% during water/LS filling and transfer operations. Redundancy measures such as modular PLCs, one-operational-one-standby pumps, and software safeguards (e.g., dual HMIs and multi-level alarms) ensure system-level fault tolerance and anti-interference capability.
These accomplishments establish a solid foundation for the safe and efficient filling of the 20-kiloton central detector, directly enabling JUNO’s primary scientific missions, including the determination of neutrino mass ordering and high-precision oscillation measurements. The system’s ability to maintain long-term LS stability and purity is essential throughout the experiment’s projected 20-year lifespan.
Looking ahead, upgrade paths are under consideration to further enhance system performance, such as the implementation of model predictive control (MPC) for process optimization and machine learning–based diagnostic tools for predictive maintenance. Moreover, the FOC system provides a ready-made infrastructure for future physics programs at JUNO, such as neutrinoless double-beta decay searches, which demand extreme LS purity and stability supported by online circulation.
Beyond JUNO, the design principles presented here—including the robust integration of industrial PLCs with experimental control frameworks, hierarchical safety interlocks, and large-volume liquid handling solutions—offer valuable insights for next-generation large-scale detectors, such as Hyper-Kamiokande and DUNE, which face similar liquid-target control challenges.